dinonuggets12:

Guys i need help!!! The distributive property states that if a, b, and c are numbers or expressions, then a(b + c) = a • b + a • c. 6(5 + 6) = 6 • 5 + 6 • 6 6(5 + 6) = 30 + 36 6(5 + 6) = 66

Vocaloid:

Are you having trouble understanding the math? let's look at a(b + c) = a • b + a • c. notice how they take the variable outside the parentheses (a) and multiply it by the things inside the parentheses to get a*b, as well as a*c. finally, they add a*b + a*c. this process is called the distributive property.

dinonuggets12:

Ok*

dinonuggets12:

The commutative property of addition says you can add numbers in any order and get the same answer. Is the same true for subtraction? 22 – 12 10 12 – 22 –10 10 ≠ –10 22 – 12 ≠ 12 – 22 a – b ≠ b – a Changing the order in which you subtract numbers does change the result, so the commutative property does not work for subtraction.

Vocaloid:

so now let's apply it 6(5 + 6) take what's outside the parentheses (6) multiply it by the things inside the parentheses (6*5 and 6*6) then add (6*5 + 6*6) finally, simplify (30 + 36 = 66)

dinonuggets12:

i got the other one i need help with this one ;-;

dinonuggets12:

2 – 12 = 12 – 2 B. (3 – 11) – 5 = 3 – (11 – 5) C. 8 + 7 = 7 + 8 D. (6 + 9) + 4 = 6 + (9 + 4)

Vocaloid:

Second post: like the post says, addition is commutative. this means you can change the order of the things being added. so for 1 + 2, you can switch 1 and 2 and get the same result 1 + 2 = 3 2 + 1 = 3 also

Vocaloid:

However, with subtraction, you can't do this. 1 - 2 = -1 2 - 1 = 1 which are not the same For the problem you posted, what exactly is being asked? I didn't quite catch the question.

dinonuggets12:

lemme make another one